OpenStudy (anonymous):

A quantity is given by E *L* (V/T). where E is the permitivity of free space,L is the length, V is the potential difference ant T is a time interval.The dimensional formula for X is same as that of ?

OpenStudy (ash2326):

We know that for a parallel plate capacitor, its capacitance is given as \[C=\frac{\epsilon_o A}{D}\] Here We have \[\epsilon_o \times L\times \frac{V}{T}\] Let's multiply and divide this by L \[\epsilon_o \times \frac{L^2}{L} \times \frac{V}{T}\] \[\underline{\epsilon_o \times \frac{L^2}{L}} \times \frac{V}{T}\] This underlined term has units of capacitance So, we can write this as C, \[C\times \frac V T\] We knowcharge Q=CV so this in most simplified terms is \[\frac{Q}{T}\] @JOYAL What do you think is this?